OpenStudy (anonymous):

please help: simplify the radical expression: cube root of 81x^4 divided by cube root of 3x

OpenStudy (anonymous):

cube root of 3x^2

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

\[\sqrt[3]{\frac{81x^4}{3x}}=\sqrt[3]{27x^3}=3x\]

OpenStudy (anonymous):

as \[\frac{\sqrt[3]{81x^4}}{\sqrt[3]{3x}}=\sqrt[3]{\frac{81x^4}{3x}}\]

OpenStudy (anonymous):

its 3x

OpenStudy (anonymous):

3x is my answer

OpenStudy (anonymous):

tnx